OpenStudy (anonymous):

Simplify: x+3/9 / x+12/6 a. (x+3)(x+12) / 6 b. 2(x+3) / 3(x+12) c. 2x / x+12 d. 1/6

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

\[\huge \frac{x + \frac{3}{9}}{x + \frac{12}{6}}\] Like that?

OpenStudy (anonymous):

No :P

OpenStudy (anonymous):

9 is under x+3 on the top and 6 is under x+12 on the bottom

OpenStudy (anonymous):